# A Differentiable Model of Nucleic Acid Dynamics

### v1

@@ 4 Materials and Methods
Here we provide details on the oxDNA model as well as the method for stochastic gradient estimation given knowledge of unnormalized probabilities develoed independently by Zhang et al. [50] and Thaler and Zavadlav [51].
4.1 oxDNA
The oxDNA force field is a coarse-grained model of DNA first introduced in 2010 by Thomas Ouldridge [57]. Originally intended to capture DNA properties salient to DNA nanotechnological devices – hybridization, mechanical properties of both single- and double-stranded DNA – oxDNA has since been profitably used in a host of biophysical contexts [73, 74, 75] in addition to becoming a workhorse in the DNA nanotechnology community [40, 54, 60, 76]. The model, widely validated by experiments, has been used to study phenomenon ranging from small-scale strand displacement reactions [41, 44, 45] to the gelation and crystallization of DNA origamis [42, 43, 46, 47]. In its original form, the only sequence specificity in oxDNA was the restriction of hydrogen bonds to valid base pairs (i.e. A-T and G-C). Since its inception, oxDNA has been extended to include base-pair specific hydrogen-bonding and stacking parameters and major and minor grooving. oxDNA has also been tuned to model RNA as well as DNA-RNA interactions.
Under the oxDNA model, DNA is modelled as a string of rigid nucleotides. Each nucleotide is represented as a rigid body consisting of a backbone repulsion site, a stacking site, and a hydrogen-bonding site. All interactions are pairwise and the model considers six types of interactions: backbone connectivity, excluded volume, hydrogen bonding, stacking, cross stacking, and coaxial stacking. More specifically, the potential energy is given by
@@
where $nn$ denotes the set of consecutive bases within strands. See [57] for details. The latest version of oxDNA, oxDNA 2, contains an additional Debye-Hückel term to model electrostatics. All optimizations presented in this work use the oxDNA 2 model unless otherwise stated.
@@ Acknowledgments
We thank Petr Sulc Tom Ouldridge, Jonathan Doye, Lorenzo Rovigatti, Erik Poppleton, and Ard Louis for support and helpful discussions surrounding the oxDNA force field and the oxDNA ecosystem. We also thank Petr Sulc and Erik Poppleton for inspiring sequence-structure optimization. We thank Salvatore Assenza and Rubén Pérez for their helpful feedback regarding the calculation of mechanical properties, including providing simulation protocols for stretch-torsion simulations in LAMMPS. We thank Jamie Smith for helpful conversations relating to stochastic gradient estimation. This material is based upon work supported by the NSF AI Institute of Dynamic Systems (#2112085) and the Office of Naval Research (N00014-17-1-3029).
@@ A.2 Differentiable Molecular Dynamics
When embedded within an automatic differentiation framework, gradients of this computation can be efficiently computed with respect to $\theta$. Assume an objective function dependent on the final state $\mathbf{x}_{N}$ and control parameters $\theta$, denoted $O(\mathbf{x}_{N},\theta)$. Given the stochastic nature of MD trajectories, the common objective is the expectation of such functions:
@@
It is important to note that (i) objective functions can also be defined to depend on the entire trajectory rather than solely on the final state, and (ii) in equilibrium systems with long simulations, states separated by sufficiently large time intervals can be interpreted as distinct trajectories under standard assumptions of ergodicity. However, for clarity, this discussion will focus on current formulations.
Since gradients must be computed while considering the simulation process, which requires repeated application of $\mathcal{S}$, computational and memory limitations may arise. To illustrate, consider the gradient in Equation 22 for one-dimensional space (see Ref. [48] for details):
@@ A.3 Implicit Differentiation
An alternative approach to differentiate through optimization procedures is to implicitly relate the solution of an optimization problem to its inputs using optimality conditions. Consider an optimization procedure $y:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that maps an input $x\in\mathbb{R}^{n}$ to a solution $y(x)\in\mathbb{R}^{m}$ that satisfies an optimality condition described by the relation $R(y(x),x)=0$. When the relation $R$ is continuously differentiable and $\partial R_{y}$ evaluated at $(y(x),x)$ is square invertible, the implicit function theorem provides expression for the implicit relationship between $x$ and $y(x)$:
@@
Historically, this approach required complicated, case-specific derivations that made impliicit differentiation difficult to use in practice. However, Blondel et al. [62] developed a method for automatic implicit differentiation that uses automatic differentiation to compute $\partial_{x}y(x)$ without deriving $\partial R_{x}(y(x),x)$ and $\partial R_{y}(y(x),x)$ by hand. We use this method in this work to compute derivatives of line fitting procedures involved in the calculation of thermodynamic and mechanical properties of DNA.
Appendix B Notation
@@
In oxDNA, a nucleotide has three interaction sites used in the energy calculation – a backbone repulsion site, a stacking site, and a hydrogen bonding site. For a particular nucleotide with centre of mass $\mathbf{r}=\mathbf{R_{i}}\in\mathbb{R}^{3}$, the locations of these interaction sites in the space-fixed frame are given by:
@@ Appendix C Benchmarking
We first sought to benchmark performance for forward simulations via a pure JAX implementation of oxDNA vs. the standalone implementation. Our pure JAX implementation involves an implementation of the oxDNA energy function in JAX and relies on JAX-MD for integration and force calculation via automatic differentiation. We consider a small system (8 base pairs) and a large system (500 base pairs) on CPU and GPU, respectively; in Ref. [40], Sengar et al. highlight that small systems do not benefit from GPU acceleration. We also employ neighbor lists to accelerate the force calculation for the 500 base pair system. We observe a difference in runtime commensurate with the statistics reported by Schoenholz and Cubuk in their original presentation of JAX-MD [53]. For the large system, we observe a 64% increase in runtime. We observe a significantly larger runtime cost to simulate the small system in JAX-MD due to the fixed overhead of JIT-compilation.
[…]
We next sought to benchmark the overhead of gradient calculation via our JAX implementation. We considered only the 8 base pair system on both CPU and GPU. We also compared vanilla backpropagation through the unrolled trajectory with checkpointed gradients (every 100 timesteps), a method which can enable gradient calculation through longer simulations by reducing memory at the cost of runtime. For this system, vanilla backpropagation imposes a significant runtime overhead compared to forward simulations (approximately 450% and 850% on CPU and GPU, respectively). Gradient checkpointing imposes an additional 7-9% runtime increase. Note that this cost may be larger for longer time scales or more complex systems, but this is impossible to evaluate empirically as gradients cannot be computed for such simulations without checkpointing.
@@ D.1.1 Definition
The pitch of a double-stranded DNA (or RNA) duplex describes the height of one complete helix turn, measured parallel to the helical axis; this is typically reported as the number of base pairs per helical turn. Consider a duplex comprised of two single strands, $\alpha$ and $\beta$, each with $n$ nucleotides. This duplex contains $n$ base pairs where the $i^{th}$ base pair involves the pairing of the $i^{th}$ nucleotide of $\alpha$ and the $(n-i)^{th}$ nucleotide of $\beta$. For two adjacent base pairs $i$ and $j$, we define the local helix axis as
@@
In practice, the first two base pairs at each end of the duplex are not included in the calculation to remove the effects of fraying. We can then define the (differentiable) average $\theta$ across many equilibrium states
@@
Figure S3: Optimization of oxDNA parameters for target propeller twist. A. The loss values over time, defined as the RMSE between the calculated propeller twist and a target propeller twist of 19°. B. The five most significant parameters ranked by the absolute value of their gradients.
@@ D.1.3 Optimization via Score-Function Estimator
We then sought to evaluate the performance of the score function estimator). This requires obtaining a differentiable expression for the probability of each trajectory. The Langevin integrator in JAX-MD implements the integration scheme of Davidchak et al. [79] that follows a “BAOAB” splitting scheme [80], one choice for splitting the Langevin equation into deterministic and stochastic components. We represent the probability of a given timestep as the probability of the stochastic step in the integration scheme, and therefore the probability of a trajectory as the product of the probabilities over all timesteps. We extract this probability from JAX-MD and it is differentiable via JAX. All simulations were initialized with states obtained from $1,000$ equilibration steps and all optimizations used a learning rate of $0.001$.
Given a simulation length and batch size, we use this probability to obtain gradient estimates for pitch optimization for optimization. We evaluate a range of simulation lengths and batch sizes with a bias towards shorter simulations and higher batch sizes given the high variance introduced by longer simulations. In practice, we do not find any hyperparameter set that yields optimized parameter values. See Figure S2 for our results.
@@ D.2 Propeller Twist
Again, consider a duplex comprised of two single strands, $\alpha$ and $\beta$, each with $n$ nucleotides where the $i^{th}$ base pair involves the pairing of the $i^{th}$ nucleotide of $\alpha$ and the $(n-i)^{th}$ nucleotide of $\beta$. For the $i^{th}$ base pair, let $\mathbf{a}_{2,i}^{\alpha}$ and $\mathbf{a}_{2,n-i}^{\beta}$ denote the vectors normal to the plane of the base for each of two base paired nucleotides (see Appendix B for the definition of $\mathbf{a}_{2}$). The propeller twist of the $i^{th}$ base pair is
@@ E.1.1 Passive calculation
Crucially, the entire calculation of $L_{ps}$ given a set of reference states can be made differentiable by combining DiffTRE and implicit differentiation. $\mathbb{E}[O_{m}(\mathbf{R})]$ is differentiable for all $m$ via reweighting and implicit differentiation can be used to differentiate the line-fitting procedure described in Equations 51 and 52 in the case where an analytic solution is not available.
@@
Figure S5: Optimization of stretch modulus via a worm-like chain fit in which the persistence length is constrained via a passive calculation. A. The stretch modulus and corresponding persistence length over time. B. The force extension curves corresponding to the initial and optimized parameters.

### v2

@@ 4 Materials and Methods
Here we provide details on the DNA, RNA, and DNA-protein models optimized in this work as well as the method for stochastic gradient estimation given knowledge of unnormalized probabilities developed independently by Zhang et al. [30] and Thaler and Zavadlav [29].
4.1 Force Fields
4.1.1 oxDNA and oxRNA
The oxDNA force field is a coarse-grained model of DNA first introduced in 2010 by Thomas Ouldridge [69]. Originally intended to capture DNA properties salient to DNA nanotechnological devices – hybridization, mechanical properties of both single- and double-stranded DNA – oxDNA has since been profitably used in a host of biophysical contexts [102, 103, 104] in addition to becoming a workhorse in the DNA nanotechnology community [60, 65, 72, 105]. The model, widely validated by experiments, has been used to study phenomena ranging from small-scale strand displacement reactions [37, 106, 107] to the gelation and crystallization of DNA origami [61, 62, 108, 109]. In its original form, the only sequence specificity in oxDNA was the restriction of hydrogen bonds to valid base pairs (i.e. A-T and G-C). Since its inception, oxDNA has been extended to include base-pair specific hydrogen-bonding and stacking parameters [23] and major and minor grooving [26].
In the oxDNA model, DNA is modelled as a string of rigid nucleotides. Each nucleotide is represented as a rigid body consisting of a backbone repulsion site, a stacking site, and a hydrogen-bonding site. All interactions are pairwise and the model considers six types of interactions: backbone connectivity, excluded volume, hydrogen bonding, stacking, cross stacking, and coaxial stacking. More specifically, the potential energy is given by
@@ 4.1 Force Fields
where $nn$ denotes the set of consecutive bases within strands; see [69] for details. The latest version of oxDNA, oxDNA2, contains an additional Debye-Hückel term to model electrostatics. All optimizations presented in this work use the oxDNA2 model unless otherwise stated.
@@
[…]
4.1.2 DNA-Protein Model
In Ref. [25], Procyk et al. introduce a coarse grained model of DNA-protein hybrids. The combined model includes energetic descriptions of (i) DNA-DNA, (ii) protein-protein, and (iii) DNA-protein interactions. DNA-DNA interactions are described by the oxDNA model. Protein-protein interactions are given by an Anisotropic Network Model (ANM), which represents a protein with a known structure as beads connected by spring. Each residue is modelled as a single bead. To capture fluctuations of the protein backbone, spring constants are chosen to best fit B-factors of the $\alpha-\text{carbons}$ given in the original PDB structure. This is performed automatically via oxView [95]. DNA-protein interactions are simply modelled as excluded volume effects, neglecting any attractive or non-steric repulsive interactions. Procyk additionally develop more complex protein-protein interactions that include torsional and bending potentials, though we focus on the ANM model in this work.
We extend this model by parameterizing DNA-protein interactions via a sequence-specific Morse potential. Specifically, given an interacting pair consisting of a DNA nucleotide of type $i$ (corresponding to one of four possible nucleotides) and an amino acid of type $j$ (corresponding to one of 20 possible amino acids), we introduce an attractive interaction:
$V_{\text{morse}}^{ij}=\epsilon^{ij}\left(1-\exp\left(-\alpha^{ij}\left(r-r^{ij}_{0}\right)\right)\right)^{2}$ (6)
where $r$ is the inter-particle distance, $r^{ij}_{0}$ is the equilibrium bond length, $\epsilon^{ij}$ is the sequence-specific strength, and $\alpha^{ij}$ is the sequence-specific width. Thus, this new interaction is parameterized via $\boldsymbol{\epsilon},\boldsymbol{\alpha},\mathbf{r_{0}}\in\mathbb{R}^{4\times 20}$. In practice, we parameterize individual Morse potentials for both the backbone site and hydrogen bonding site and therefore fit six $4\times 20$ matrices (three for each site).
@@ 4.2 Differentiable Monte Carlo (DiffTRE)
4.3 Conflict-Free Gradients
Here we generically describe conflict-free gradient methods, and subsequently the ConFIG method of Liu et al. employed in this work [59]. We largely follow the notation and presentation of Ref. [59].
Consider an optimization problem in which the goal is to find $\theta$ that simultaneously minimizes $m$ individual loss functions, $\{\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{m}\}$ where the total loss function is expressed as
$\mathcal{L}(\theta)=\sum_{i=1}^{m}\mathcal{L}_{i}(\theta)$ (13)
Naively, direct differentiation of this objective yields
$\nabla_{\theta}\mathcal{L}(\theta)=\sum_{i=1}^{m}\nabla_{\theta}\mathcal{L}_{i}(\theta)$ (14)
However, in such problems, it is typical that optimization becomes stuck in a local minimum of a specific loss term due to the conflict between losses. Consider a gradient update $\mathbf{g}^{\text{total}}$, e.g. $\mathbf{g}^{\text{total}}=\nabla_{\theta}\mathcal{L}(\theta)$ as in Equation 14. A gradient update will conflict with the decrease of $\mathcal{L}_{i}$ if $\mathbf{g}_{i}^{\top}\mathbf{g}^{\text{total}}<0$ where $\mathbf{g}_{i}=\nabla_{\theta}\mathcal{L}_{i}(\theta)$ [59, 113, 114]. One solution is to construct $\mathbf{g}^{\text{total}}$ such that
$\mathbf{g}^{\text{total}}=[\mathbf{g}_{i},\mathbf{g}_{2},\ldots,\mathbf{g}_{m}]^{-\top}\mathbf{w}$ (15)
where $w_{i}>0$ for all $i$ and $M^{-\top}$ is the pseudoinverse of the transposed matrix $M^{\top}$.
In this work, we apply the recently-developed Conflict-Free Inverse Gradients (ConFIG) method of Liu et al. [56]. In this scheme, $\mathbf{g}^{\text{total}}$ is constructed as follows:
$\mathbf{g}^{\text{total}}$ $=\left(\sum_{i=1}^{m}\mathbf{g}_{i}^{\top}\mathbf{g}_{u}\right)\mathbf{g}_{u}$ (16)
$\mathbf{g}_{u}$ $=\mathcal{U}\left[\left[\mathcal{U}(\mathbf{g}_{1}),\mathcal{U}(\mathbf{g}_{2}),\ldots,\mathcal{U}(\mathbf{g}_{m})\right]^{-\top}\mathbf{1}_{m}\right]$ (17)
where $\mathcal{U}(\mathbf{g}_{i})$ is a normalization operator and $\mathbf{1}_{m}$ is a unit vector with $m$ components that ensures a uniform rate of decrease across all loss terms. This unit vector can be replaced with custom direction weights $\hat{\mathbf{w}}$ to control the relative importance of individual loss terms, but we set $\hat{\mathbf{w}}=\mathbf{1}_{m}$ in this work. See Ref. [56] for a complete derivation.
@@ Acknowledgments
We thank Petr Sulc, Tom Ouldridge, Jonathan Doye, Lorenzo Rovigatti, Erik Poppleton, and Ard Louis for support and helpful discussions surrounding the oxDNA force field and the oxDNA ecosystem. We also thank Petr Sulc and Erik Poppleton for inspiring sequence-structure optimization. We thank Salvatore Assenza and Rubén Pérez for their helpful feedback regarding the calculation of mechanical properties, including providing simulation protocols for stretch-torsion simulations in LAMMPS. We thank Jamie Smith for helpful conversations relating to stochastic gradient estimation. R.K.K. thanks Stefania Ketzetzi for helpful feedback on the manuscript. This material is based upon work supported by the NSF AI Institute of Dynamic Systems (#2112085) and the Office of Naval Research (N00014-17-1-3029).
@@ A.2 Differentiable Molecular Dynamics
Figure S1: Traditional means of stochastic gradient estimation for differentiable molecular dynamics (MD). A. Under the pathwise estimator, gradients are directly calculated via a differentiable simulator. B. Alternatively, the score function estimator scales the gradients of (log) probabilities of individual trajectories by a reward defined for each trajectory. Differentiable (MD) typically relies on the pathwise estimator, though Engel et al. also used the score function estimator [115].
When embedded within an automatic differentiation framework, gradients of this computation can be efficiently computed with respect to $\theta$. Assume an observable dependent on the final state $\mathbf{x}_{N}$ and control parameters $\theta$, denoted $O(\mathbf{x}_{N},\theta)$. Given the stochastic nature of MD trajectories, the common objective is the expectation of such functions:
@@
It is important to note that (i) objective functions can also be defined to depend on the entire trajectory rather than solely on the final state, and (ii) in equilibrium systems with long simulations, states separated by sufficiently large time intervals can be interpreted as distinct trajectories under standard assumptions of ergodicity. However, for clarity, this discussion will focus on the above formulation. Figure S1 depicts the methods for stochastic gradient estimation described above in the context of differentiable MD.
Since gradients must be computed through the simulation process, which requires repeated application of $\mathcal{S}$, computational and memory limitations may arise. To illustrate this, consider the gradient in Equation 27 for one-dimensional space (see Ref. [31] for details):
@@ A.3 Implicit Differentiation
An alternative approach to differentiating through optimization procedures is to implicitly relate the solution of an optimization problem to its inputs using optimality conditions. Consider an optimization procedure $y:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that maps an input $x\in\mathbb{R}^{n}$ to a solution $y(x)\in\mathbb{R}^{m}$ that satisfies an optimality condition described by the relation $R(y(x),x)=0$. When the relation $R$ is continuously differentiable and $\partial R_{y}$ evaluated at $(y(x),x)$ is square invertible, the implicit function theorem provides expression for the implicit relationship between $x$ and $y(x)$:
@@
Historically, this approach required complicated, case-specific derivations that made implicit differentiation difficult to use in practice. However, Blondel et al. [68] developed a method for automatic implicit differentiation that uses automatic differentiation to compute $\partial_{x}y(x)$ without deriving $\partial R_{x}(y(x),x)$ and $\partial R_{y}(y(x),x)$ by hand. We use this method in our work to compute derivatives of line fitting procedures involved in the calculation of thermodynamic and mechanical properties of DNA.
[…]
Appendix B Model Details
@@
In oxDNA, a nucleotide has three interaction sites used in the energy calculation: a backbone repulsion site, a stacking site, and a hydrogen bonding site. For a particular nucleotide with centre of mass $\mathbf{r}=\mathbf{R_{i}}\in\mathbb{R}^{3}$, the locations of these interaction sites in the space-fixed frame are given by:
@@ Appendix C Benchmarking
We first sought to benchmark performance for forward simulations via a pure JAX implementation of oxDNA vs. the standalone implementation. Our pure JAX version involves an implementation of the oxDNA energy function in JAX and relies on JAX-MD for integration and force calculation via automatic differentiation. We consider a small system (8 base pairs) and a large system (500 base pairs) on CPU and GPU, respectively; in Ref. [60], Sengar et al. highlight that small systems do not benefit from GPU acceleration. We also employ neighbor lists to accelerate the force calculation for the 500 base pair system. We observe a difference in runtime commensurate with the statistics reported by Schoenholz and Cubuk in their original presentation of JAX-MD [36]. For the large system, simulated with GPU, we observe a 64% increase in runtime relative to the standalone oxDNA CUDA code (see Fig. S2A). We observe a significantly larger runtime cost to simulate the small system in JAX-MD due to the fixed overhead of JIT-compilation.
We next sought to benchmark the overhead of gradient calculation via our JAX implementation. We considered only the 8 base pair system on both CPU and GPU. We also compared vanilla backpropagation through the unrolled trajectory with checkpointed gradients (every 100 timesteps), a method which can enable gradient calculation through longer simulations by reducing memory at the cost of runtime. For this system, vanilla backpropagation imposes a significant runtime overhead compared to forward simulations (approximately 450% and 850% on CPU and GPU, respectively); see Fig. S2B. Gradient checkpointing imposes an additional 7-9% runtime increase. Note that this cost may be larger for longer time scales or more complex systems, but this is impossible to evaluate empirically as gradients cannot be computed for such simulations without checkpointing.
@@ D.1.1 Definition
The pitch of a double-stranded DNA (or RNA) duplex describes the height of one complete helix turn, measured parallel to the helical axis; this is typically reported as the number of base pairs per helical turn. Consider a duplex comprised of two single strands, each with $n$ nucleotides. This duplex contains $n$ base pairs where the $i^{th}$ base pair involves the pairing of the $i^{th}$ nucleotide strand 1 and the $(n-i)^{th}$ nucleotide of strand 2. For two adjacent base pairs $i$ and $j$, we define the local helix axis as
@@
where $\mathbf{R}$ denotes a single configuration. In practice, the two base pairs nearest each end of the duplex are not included in the calculation to remove the effects of fraying. We can then define the (differentiable) average $\theta$ across many equilibrium states
@@
Figure S4: Fitting oxDNA parameters to a target propeller twist via trajectory reweighting. A. The loss values over time, defined as the RMSE between the calculated propeller twist and a target propeller twist of 19°. Iterations in which reference state are resampled are represented as scatter points. B. The five most significant parameters ranked by the absolute value of their gradients.
@@ D.1.3 Optimization via Score-Function Estimator
We then sought to evaluate the performance of the score function estimator. This requires obtaining a differentiable expression for the probability of each trajectory. The Langevin integrator in JAX-MD implements the integration scheme of Davidchak et al. [116] that follows a “BAOAB” splitting scheme [117], one choice for splitting the Langevin equation into deterministic and stochastic components. We represent the probability of a given timestep as the probability of the stochastic step in the integration scheme, and therefore the probability of a trajectory as the product of the probabilities over all timesteps. We extract this probability from JAX-MD and it is differentiable via JAX. All simulations were initialized with states obtained from $1,000$ equilibration steps and all optimizations used a learning rate of $0.001$.
Given a simulation length and batch size, we use this probability to obtain gradient estimates for pitch optimization for optimization. We evaluate a range of simulation lengths and batch sizes with a bias towards shorter simulations and higher batch sizes given the high variance introduced by longer simulations. In practice, we do not find any hyperparameter set that yields optimized parameter values. See Figure S3B for our results.
@@ D.2 Propeller Twist
Again, consider a duplex comprised of two single strands, $\alpha$ and $\beta$, each with $n$ nucleotides where the $i^{th}$ base pair involves the pairing of the $i^{th}$ nucleotide of $\alpha$ and the $(n-i)^{th}$ nucleotide of $\beta$. For the $i^{th}$ base pair, let $\mathbf{a}_{2,i}^{\alpha}$ and $\mathbf{a}_{2,n-i}^{\beta}$ denote the vectors normal to the plane of the base for each of two base paired nucleotides (see above for the definition of $\mathbf{a}_{2}$). The propeller twist of the $i^{th}$ base pair is
@@ D.3 Sequence-Structure Pair
Figure S6: Optimization of stretch modulus via a worm-like chain fit in which the persistence length is constrained via a passive calculation. A. The stretch modulus and corresponding persistence length over time. B. The force extension curves corresponding to the initial and optimized parameters. Iterations in which reference state are resampled are represented as scatter points.
@@ E.1.1 Passive calculation
Crucially, the entire calculation of $L_{ps}$ given a set of reference states can be made differentiable by combining DiffTRE and implicit differentiation. $\mathbb{E}[O_{m}(\mathbf{R})]$ is differentiable for all $m$ via reweighting and implicit differentiation can be used to differentiate the line-fitting procedure described in Equations 56 and 57 in the case where an analytic solution is not available.